How do you find what the actual yearly interest rate is if you buy something for $320.00 and you pay $20.00 down and $34.00 a month for 10 months.
Answer by richwmiller(9144)

let's find out how how much was paid.
c=20+34*10=360
now what was the interest
360-320=40
now the annual percentage
i=p*r*t
40=320*r*10/12=3/20=1.5/10=15/100=15%
regular percentage would be 40/320=1/8=12.5 %